OpenStudy (sedatefrog712):

I'm having some trouble figuring this question out A company distributes free candies to all the students of x schools. Each school has (x + 1) classes. The number of students in each class is 3 more than the number of classes in each school. Each student is given 4 candies. Part A: Write an expression to show the total number of candies distributed by the company in x schools.

OpenStudy (sedatefrog712):

Part B: What would x(x + 1) represent? When simplified, what would be the degree and classification of this expression? How can you calculate the total number of students in each school? @AkashdeepDeb @ganeshie

OpenStudy (sedatefrog712):

@SolomonZelman

OpenStudy (solomonzelman):

You are learning about a sigma notation right now?

OpenStudy (sedatefrog712):

Yes sir

OpenStudy (solomonzelman):

Okay, so you know that a1 is x+1, and that d=3 and in terms of candy, a1 = 4(x+1) where a2=4(x+4), a3=(x+7), a4=(x+10) and on.... With X terms.

OpenStudy (solomonzelman):

I mean a3=4(x+7) and a4=4(x+10)

OpenStudy (solomonzelman):

and the x is unknown

OpenStudy (solomonzelman):

\[\sum_{n=1}^{x}~4(~3n+x)\]

OpenStudy (solomonzelman):

So in terms of candy\[\sum_{n=1}^{x}~4(~3n+x)\] and in terms of student\[\sum_{n=1}^{x}~~3n+x\]

OpenStudy (sedatefrog712):

ok

OpenStudy (sedatefrog712):

would I distribute any thing

OpenStudy (solomonzelman):

why distribute ?

OpenStudy (sedatefrog712):

because it says Write an expression to show the total number of candies distributed by the company in x schools.

OpenStudy (solomonzelman):

Well, I think I did it when I wrote the Σ in terms of candy.

OpenStudy (sedatefrog712):

o ok I see I see thank you I feel like an idiot

OpenStudy (solomonzelman):

And would YOU distribute anything, is a different question. If you were them you would probably not :P